OpenStudy (jkbo):

Choose a counterexample that proves that the conjecture below is false. Other than the number 1, there are no numbers less than 100 that are both perfect squares and perfect cubes. (1 point) 36 64 16 8

OpenStudy (btaylor):

64=8^2=4^3

OpenStudy (jkbo):

64?

OpenStudy (btaylor):

yes.

OpenStudy (jkbo):

k thanks